If you walk chisale to st Francis , a distance of 6km in50minutes, find the average speed in m/s


1
Expert's answer
2021-02-23T10:03:11-0500
"v=\dfrac{s}{t}=\dfrac{6\ km\cdot\dfrac{1000\ m}{1\ km}}{50\ min\cdot\dfrac{60\ s}{1\ min}}=2\ \dfrac{m}{s}."